OpenStudy (anonymous):

divide and simplify to lowest term. 2x^2-6x/3x-2 divided by x-3/3x-2 /= over

OpenStudy (anonymous):

Care to simplify the question?

OpenStudy (anonymous):

or wait, I think I got it. another name for division is inverse multiplication.

OpenStudy (anonymous):

So, \[\frac{ 2x^{2}-6x }{ 3x+2 } \div \frac{ x-3 }{ 3x+2 }\] is actually \[\frac{ 2x^{2}-6x }{ 3x+2 } \times \frac{ 3x+2 }{ x-3 }\]

OpenStudy (anonymous):

Which ultimately gives you 2x.

OpenStudy (anonymous):

 how does that turn into 2x?

OpenStudy (anonymous):

3x-2 in both places. :\

OpenStudy (anonymous):

the +2 bit was a mistake (refer to the question asked)

OpenStudy (anonymous):

oh ok. oops lol. thanks soo much :)

OpenStudy (anonymous):

Glad to be of use. ;)